### v1

@@ Abstract.
In the evolving e-commerce field, recommendation systems crucially shape user experience and engagement. The rise of Consumer-to-Consumer (C2C) recommendation systems, noted for their flexibility and ease of access for customer vendors, marks a significant trend. However, the academic focus remains largely on Business-to-Consumer (B2C) models, leaving a gap filled by the limited C2C recommendation datasets that lack in item attributes, user diversity, and scale. The intricacy of C2C recommendation systems is further accentuated by the dual roles users assume as both sellers and buyers, introducing a spectrum of less uniform and varied inputs. Addressing this, we introduce MerRec, the first large-scale dataset specifically for C2C recommendations, sourced from the Mercari e-commerce platform, covering millions of users and products over 6 months in 2023. MerRec not only includes standard features such as user_id, item_id, and session_id, but also unique elements like timestamped action types, product taxonomy, and textual product attributes, offering a comprehensive dataset for research. This dataset, extensively evaluated across three recommendation tasks, establishes a new benchmark for the development of advanced recommendation algorithms in real-world scenarios, bridging the gap between academia and industry and propelling the study of C2C recommendations.
@@
Figure 1. Comparison of B2C and C2C E-commerce Platforms: This illustration highlights the differences in product descriptions between B2C and C2C platforms. B2C platforms typically feature consistent and professionally crafted merchandise descriptions. Conversely, product descriptions on C2C platforms are often more varied and less standardized.
@@ 1. Introduction
In the era of flourishing e-commerce, recommendation systems have emerged as pivotal tools in shaping user experience and driving customer engagement. These systems are instrumental in navigating the vast array of products and services offered online, making them a cornerstone of modern e-commerce platforms (Isinkaye et al., 2015; Zhao et al., 2021; Fan et al., 2022; Zhang et al., 2019). Currently, the literature of recommendation systems has been dominated by Business-to-Consumer (B2C) models (Alamdari et al., 2020), which have been extensively studied and refined over the years. These models typically revolve around businesses recommending products to individual consumers, leveraging structured data and well-defined user behavior patterns (Li and Karahanna, 2015). More recently, with the advent of digital marketplaces and the shift towards more inclusive and flexible e-commerce platforms, the Consumer-to-Consumer (C2C) (Ovaska et al., 2024) model has developed prosperously. In C2C e-commerce, individuals are both the sellers and buyers of products, creating a dynamic and multifaceted marketplace. This shift marks a significant trend for E-commerce, highlighting the need for recommendation systems that cater to these unique environments. Unlike their B2C counterparts, C2C platforms present unique challenges due to their unstructured nature, the dual roles of users, and the potential absence of conventional identifiers like Stock Keeping Units (SKUs).
@@
The comprehensiveness makes MerRec a multipurpose dataset and can be utilized for diverse tasks related to recommendations. Here we focus on three different tasks: ❶ Click-through rate (CTR) prediction, which is a classical task where a recommendation model learns to anticipate whether a user click is likely to happen given some context. ❷ Session-based Recommendation (SBR), or sequential recommendation, which is a widely used modeling technique to predict the next item, given a list of interacted items. ❸ Multi-task Learning for Recommendation (MLR), where the recommendation model is expected to predict different types of user actions such as item_view and item_like. Our benchmarks include both sophisticated baseline models and straightforward heuristic approaches, providing a comprehensive overview of MerRec’s utility.
@@ 2.1. E-commerce Recommendation Systems
The emergence of Consumer-to-Consumer (C2C) platforms like eBay, Etsy, and Mercari has significantly altered the dynamics of the e-commerce ecosystem. These platforms empower individuals to both offer and acquire goods directly, fostering a marketplace characterized by its vast diversity yet complicated by intricate user interactions and varied transaction models. Research into C2C platforms has delved into the socio-economic behaviors of users, the mechanisms of trust and reputation systems, and the distinct challenges posed by a marketplace driven by its participants (Cano et al., 2023). Despite this, crafting recommendation systems for such environments is fraught with challenges. The lack of uniform product identifiers (SKUs), the dynamic nature of product listings, and the amalgamated roles of users as both vendors and consumers add layers of complexity. Crucially, the absence of a comprehensive, large-scale dataset presents a significant obstacle for researchers in the field, hindering the advancement of tailored recommendation solutions.
@@ 2.2. Datasets for Recommendation Systems
The availability of robust and comprehensive datasets is crucial for advancing recommendation system research. While there are well-established datasets for B2C scenarios, such as the Amazon product review dataset (Ni et al., 2019; Jin et al., 2023) and the Netflix Prize dataset (Netflix, 2020), the provision of datasets tailored to C2C recommendation systems is limited. Previous works have either adapted existing B2C datasets to simulate C2C environments or have released only limited portions of the markets without capturing the full scale and diversity of mainstream C2C platforms (Ma et al., 2018a).
@@ 3. The [Uncaptioned image] Dataset
Mercari, an online Consumer-to-Consumer (C2C) marketplace, facilitates a unique environment where users can seamlessly transition between the roles of buyer and seller, minimizing logistical hurdles. This research is primarily focused on analyzing buyer-side preferences to enhance the relevancy of recommendation tasks within this platform. From a seller’s perspective, Mercari requires the provision of detailed information for each listing, which may include individual items or bundles. Essential details such as titles, brands, categories, images, shipping cost responsibilities, price, and condition are mandated, while size and color are optional. These details are dynamic, allowing sellers to update their listings, which introduces variability in how items are represented over time.
@@
The MerRec dataset is meticulously designed to capture the intricate and evolving dynamics of user behavior and item characteristics on Mercari [^2] MerRec is an anonymized subset of Mercari data and thus cannot not be treated as a precise or comprehensive representation of production.. It aims to establish a comprehensive understanding of the relationships between users and items by analyzing user actions in conjunction with item details. This approach seeks to address the challenges posed by the C2C business model, particularly the variability of item descriptions and the absence of standard identifiers, to enhance the effectiveness of recommendation systems in such a unique marketplace environment.
@@ 3.2. Data Cleaning and Processing
    SKU Alternative Exploration: An innovative approach was adopted to explore alternatives to traditional Stock Keeping Units (SKUs) in the C2C context. A synthetic field named as product_id was created by merging brand and the most detailed category ID available, offering a novel way to study item identification in the absence of SKUs. However, it’s important to acknowledge the limitations of this approach, as it may not provide the granularity required for effective recommendation in a live marketplace setting.
The dataset is organized into six directories, each containing Parquet files for a specific month in 2023333Available soon at https://huggingface.co/datasets/mercari-us/merrec. Each file captures individual user actions, with each row detailing a unique interaction. Actions within a user’s sequence can be reconstructed by aggregating data based on user and sequence identifiers and ordered by timestamp.
@@ 4. Experiment & Analysis
In this section, we present the application of various machine learning and recommendation models to specific tasks using the MerRec dataset, demonstrating their effectiveness and performance outcomes. We refer to a previous work (Yuan et al., 2022) for the implementations of most of these models, with necessary modifications [^4] https://github.com/mercari/mercari-ml-merrec-pub-us to accommodate the unique characteristics of the MerRec dataset.
@@ 4.1. Click-Through Rate (CTR) Prediction
To limit the total amount of computation time and resource, we chose to omit comprehensive hyperparameter search and reduce the sample used specifically for the benchmark executions. We use the first among six months of data in MerRec. We set the shortest input history window sequence in this execution to be 7, where we generate the binary prediction on the 8th item. Sequences equal to 8 events become one snapshot row and those longer than 8 events become multiple snapshot rows using the rolling window. This implies we did not pad sequences shorter than 8 events (which does not exist in MerRec) to become longer sequences unlike TenRec (Yuan et al., 2022). Due to time constraints we did not specifically study the positive to negative sampling ratio on MerRec and how it affects training quality of the benchmark models, thus no negative sampling is enforced in this work and is left to the reader to explore.
@@ 4.2. Session-based Recommendation (SBR)
For Mercari’s C2C marketplace, each item is a single quantity listing, where it can be purchased only once and it needs to be removed from the predictions after being purchased. Additionally, with a large number of item listing created daily, the output space of the recommendation model will need to be updated constantly. For these reasons, existing language model based sequential recommenders are ill-suited for a dynamic C2C marketplace like Mercari.
@@
To address the aforementioned challenges of a C2C marketplace, we introduce a Three-Tower Transformer based architecture called Mercatran, inspired by the Two-Tower architecture proposed by Covington et al. (Covington et al., 2016) for YouTube recommendations, to generate product/item id independent recommendations. Mercatran is designed to capture long user history and generate diverse and multi-time step recommendations for a dynamic marketplace like Mercari.
@@
Unlike the model proposed by Covington et al. (Covington et al., 2016) where items (videos) are represented by unique id’s, in Mercatran, we encode items with the content features only, similar to Wang et al. (Wang et al., 2021a). The output of the model is modified to produce trained embeddings directly instead of tokens (corresponding to words in a language model), which is then resolved to items using vector search. Both Covington et al. (Covington et al., 2016) and Wang et al. (Wang et al., 2021a) use a two-tower architecture to encode users and items, however, we found this architecture to be lacking in diversity where the recommendations generated are similar to the last interacted item. The model is unable to learn from older interactions even when using GRUs for the user tower (Wang et al., 2021a), since the output is a single embedding which can’t capture diverse user history. We built a solution conceptually similar to this called Mercatran V1 as the precursor of our current implementation. While it was very successful in driving orders, conversion and views (see Table 5), we observed lack of diversity, where the model would only exploit the most recent interaction to recommend items.
@@
To capture longer user-history and generate multi-time step recommendations, we propose a Three-Tower architecture called Mercatran V2, where we use Two-Towers to encodes users and a third tower to encode items. The two-towers to encode users are needed to generate recommendations into the future auto-regressively, where the first tower is a standard transformer-based encoder which can encode long user-history and the second tower is a transformer-based decoder that can generate recommendations into the future. A third tower (a transfomer-based encoder, we call item encoder) is needed to provide feedback to the user towers during training, since it’s fed the ground truth. To train the model, we optimize a symmetric cross-entropy loss over the cosine similarity scores of the predictions from the user-decoder and the associated predictions from the item-encoder, for each step, similar to Sohn (Sohn, 2016) and Radford et al. (Radford et al., 2021).
@@
During serving, the user towers are used to generate multi-step embeddings that can be used to query items in a vector database, while the third tower is used to create embeddings to index items in the vector database. The third tower can be a standard MLP. We use a transformer-based encoder without self-attention to function as the third tower.
@@
Figure 7 shows the overall architecture of our three-tower model. In the current implementation, given user-history, at maximum 22 interactions, the model predicts the next 4 interactions (as embeddings), and these embeddings are resolved to items with vector search. While this system does suffer from sampling bias (Yang et al., 2020; Yi et al., 2019) due to the training data only containing user interactions, generating multi-step recommendations introduces diversity in the results aiding in providing more visibility to a more diverse set of items.
@@
Table 4. Offline evaluation of Mercatran measured at 3 levels of granularity, item (fine), category (coarse) and brand (coarse). The steps represent the 4 forecasted steps into the future. As can be seen Mercatran maintains strong numbers even at 4 steps into the future.
@@
Pred. Type Step nDCG @5 nDCG @20 Recall @5 Recall @20 Tr./Val Time (Hrs) VRAM (GB)
Item 1 0.0407 0.0615 0.0789 0.1750 11.2 8.4

### v2

@@ Abstract.
In the evolving e-commerce field, recommendation systems crucially shape user experience and engagement. The rise of Consumer-to-Consumer (C2C) recommendation systems, noted for their flexibility and ease of access for customer vendors, marks a significant trend. However, the academic focus remains largely on Business-to-Consumer (B2C) models, leaving a gap filled by the limited C2C recommendation datasets that lack in item attributes, user diversity, and scale. The intricacy of C2C recommendation systems is further accentuated by the dual roles users assume as both sellers and buyers, introducing a spectrum of less uniform and varied inputs. Addressing this, we introduce MerRec, the first large-scale dataset specifically for C2C recommendations, sourced from the Mercari e-commerce platform, covering millions of users and products over 6 months in 2023. MerRec not only includes standard features such as user_id, item_id, and session_id, but also unique elements like timestamped action types, product taxonomy, and textual product attributes, offering a comprehensive dataset for research. This dataset, extensively evaluated across four recommendation tasks, establishes a new benchmark for the development of advanced recommendation algorithms in real-world scenarios, bridging the gap between academia and industry and propelling the study of C2C recommendations. Our experiment code is available at https://github.com/mercari/mercari-ml-merrec-pub-us and dataset at https://huggingface.co/datasets/mercari-us/merrec.
@@ 1. Introduction
In the era of flourishing e-commerce, recommendation systems have emerged as pivotal tools in shaping user experience and driving customer engagement. These systems are instrumental in navigating the vast array of products and services offered online, making them a cornerstone of modern e-commerce platforms (Isinkaye et al., 2015; Zhao et al., 2021; Fan et al., 2022; Zhang et al., 2019). Currently, the literature of recommendation systems has been primarily dominated by Business-to-Consumer (B2C) models (Alamdari et al., 2020), which have been extensively studied and refined over the years. These models typically revolve around businesses recommending professionally-sold products to general public individual consumers, leveraging structured data and well-defined user behavior patterns (Li and Karahanna, 2015).
Figure 1. Comparison of B2C and C2C E-commerce Platforms: This illustration highlights the differences in product descriptions between B2C and C2C platforms. B2C platforms typically feature consistent and professionally crafted merchandise descriptions. Conversely, product descriptions on C2C platforms are often more varied and less standardized, posing challenges to the robustness of recommendations.
In more recent years, with the advent of digital marketplaces and the shift towards more inclusive and flexible e-commerce platforms, the Consumer-to-Consumer (C2C) (Ovaska et al., 2024) business model has developed prosperously. In C2C e-commerce, individuals are both the sellers and buyers of products, creating a dynamic and multifaceted marketplace. This shift marks a significant trend for E-commerce, highlighting the need for recommendation systems that cater to these unique environments. Unlike their B2C counterparts, C2C platforms present unique challenges due to their unstructured nature, the dual roles of users, and the potential absence of conventional identifiers like Stock Keeping Units (SKUs).
@@
The comprehensiveness makes MerRec a multipurpose dataset and can be utilized for diverse tasks related to recommendations. Here we focus on four different tasks: ❶ Click-through rate (CTR) prediction, which is a classical task where a recommendation model learns to anticipate whether a user click is likely to happen given some context. ❷ Session-based Recommendation (SBR), or sequential recommendation, which is a widely used modeling technique to predict the next item, given a list of interacted items. ❸ Multi-task Learning for Recommendation (MLR), where the recommendation model is expected to predict different types of user actions such as item_view and item_like. ❹ Inference Acceleration for Recommendation (IAR), which aims to train modules to provide speedup to existing session-based recommendation models. Our benchmarks include both sophisticated baseline models and straightforward heuristic approaches, providing a comprehensive overview of MerRec’s utility as an analytical testbed.
@@ 2.1. E-commerce Recommendation Systems
The emergence of Consumer-to-Consumer (C2C) platforms like eBay [^2] https://www.ebay.com/, Etsy [^3] https://www.etsy.com/, and Mercari has significantly altered the dynamics of the e-commerce ecosystem. These platforms empower individuals to both offer and acquire goods directly, fostering a marketplace characterized by its vast diversity yet complicated by intricate user interactions and varied transaction models. Research into C2C platforms has delved into the socio-economic behaviors of users, the mechanisms of trust and reputation systems, and the distinct challenges posed by a marketplace driven by its participants (Cano et al., 2023). Despite this, crafting recommendation systems for such environments is fraught with challenges. The lack of uniform product identifiers (SKUs), the dynamic nature of product listings, and the amalgamated roles of users as both vendors and consumers add layers of complexity. Crucially, the absence of a comprehensive, large-scale dataset presents a significant obstacle for researchers in the field, hindering the advancement of tailored recommendation solutions.
@@ 2.2. Datasets for Recommendation Systems
The development of advanced recommendation systems critically hinges on the availability of robust, comprehensive datasets. In the realm of B2C (Business-to-Consumer) scenarios, several well-established datasets are frequently employed. Notable examples include the Amazon product review dataset [1, 2], the Netflix Prize dataset [3], and the MovieLens dataset [4], which have all been foundational in advancing B2C recommendation system research.
In contrast, the landscape for C2C (Consumer-to-Consumer) recommendation systems is notably underrepresented in available research resources. The specific challenges of C2C environments—such as more diverse transaction types and highly individualized user interactions—are not adequately captured by traditional B2C datasets. While some researchers have attempted to adapt B2C datasets to model C2C scenarios, these adaptations often fail to fully replicate the unique dynamics of actual C2C marketplaces [5, 6]. Moreover, datasets explicitly designed for C2C recommendation systems are scarce. The few that exist, such as those derived from smaller, niche online marketplaces [7], tend to offer only a limited view of the markets, lacking the scale and diversity seen in mainstream C2C platforms [8].
This gap in dataset availability underscores a critical need in the field: the development of large-scale, diverse datasets that are tailor-made for C2C recommendation systems. Such datasets would not only enhance the realism and applicability of research but also pave the way for innovations that are finely attuned to the nuances of consumer-driven marketplaces [9].
@@ 3. The [Uncaptioned image] Dataset
Mercari, an online Consumer-to-Consumer (C2C) marketplace, facilitates a unique environment where users can seamlessly transition between the roles of buyer and seller, minimizing logistical hurdles. This research is primarily focused on analyzing buyer-side preferences to enhance the relevancy of recommendation tasks within this platform.
From a seller’s perspective, Mercari requires the provision of detailed information for each listing, which may include individual items or bundles. Essential details such as titles, brands, categories, images, shipping cost responsibilities, price, and condition are mandated, while size and color are optional. These details are dynamic, allowing sellers to update their listings, which introduces variability in how items are represented over time.
@@
The MerRec dataset is meticulously designed to capture the intricate and evolving dynamics of user behavior and item characteristics on Mercari. It aims to establish a comprehensive understanding of the relationships between users and items by analyzing user actions in conjunction with item details. This approach seeks to address the challenges posed by the C2C business model, particularly the variability of item descriptions and the absence of standard identifiers, to enhance the effectiveness of recommendation systems in such a unique marketplace environment.
@@ 3.2. Data Cleaning and Processing
    SKU Alternative Exploration: An innovative approach was adopted to explore alternatives to traditional Stock Keeping Units (SKUs) in the C2C context. A synthetic field named as product_id was created by merging brand and the most detailed category ID available, offering a novel way to study item identification in the absence of SKUs. However, it’s important to acknowledge the limitations of this approach, as it may not provide the granularity required for effective recommendation in a live marketplace setting. See Figure 6 for an illustrative example of this constraint.
    Figure 6. Average prices of different iPhone series (red) among all items (purple) within the same brand-C2 combination (i.e. product_id) of Apple-Smartphones. Without SKU-tagging, this challenges the recommendation system to identify, differentiate and precisely promote the correct product. The price and product feature variation is high even in this example, and systems should not assume users to be broadly interested in all products under the same product_id.
The dataset is a single, monolithic table that is sharded into 6 folders, each folder corresponding to a specific month in 2023, where there are a few hundreds of parquet shards within each such folder. When loading and simply concatenating all parquet files inside one of the 6 folders, you have all the rows MerRec includes for that month. When concatenating all parquet files across all 6 folders, you then have the entire dataset [^4] While we acknowledge that this format does imply some suboptimality in terms of raw disk storage size, we ultimately chose this format acknowledging the fact that (1) this dataset is practically large enough that sharding and batched loading is generally inevitable on broadly available training hardware or setup, and (2) each row is already pre-formatted in a way that conveniently supplies all the features in a flattened way, as well as (3) the items, when showing up multiple times throughout the dataset, may look different between the instances due to our dataset being a snapshot-oriented design (meaning items are faithfully represented back in time per each instance).. Each single parquet file contains individual user actions, with each row detailing a unique interaction instance. Actions within a user’s sequence can be reconstructed by aggregating data based on user and sequence identifiers and ordered by timestamp. See Figure 7 for a structural overview.
Figure 7. Illustrative overview of the dataset organization.
@@
3.3. Comparison to Other E-commerce Datasets
We compare MerRec against other open large-scale e-commerce recommendation datasets in Table 2. To our best awareness, Retailrocket (Zykov et al., 2022) and MerRec are the only datasets in this list where item instances are back-in-time snapshots. Amongst the datasets, MerRec provides the highest number of unique items (+72% against Amazon (Hou et al., 2024)), interaction instances (+122% against Amazon (Hou et al., 2024)), number of unique interaction types (+50% against DIGINETICA [^5] https://competitions.codalab.org/competitions/11161), most amount of categories (+103.6% against Retailrocket (Zykov et al., 2022)), number of sessions (+2358% against YOOCHOOSE [^6] https://recsys.acm.org/recsys15/challenge/), most generally feature-comprehensive by far while being the most up-to-date in recency. MerRec is also the most transparent in dataset production process among these by far.
Table 2. Comparison with other large scale e-commerce datasets. Entries with explicit quoted numbers are either from claims of official work, or from our own calculation where feasible. Cells with question mark represent missing public information absent from original work or practically infeasible to gather or recover. Some columns (e.g. condition, shipment paying party) are omitted here to constrain the table size.
Dataset Market Type Users Items Interactions No. Interaction Types Categories / Leveled Brand Price Color Size Timestamp Item Tokens SKU / UPC Covered Year Sessions
MerRec (Ours) C2C (General) 5.56M 83.07M 1.27B 6 3399 / Yes Yes Yes Yes Yes Yes 18.86B No Half of 2023 227.16M
Amazon (Hou et al., 2024) B2C (General) 54.51M 48.19M 571.54M 2 33 / No Yes Yes Yes Yes Yes 30.78B Yes 1996-2023 No
Tmall (Tianchi, 2018) B2C (General) 645.37K 2.35M 44.52M 2 72 / No No No No No Yes No Yes Half of 2015 200.28K
Amazon-M2 (Jin et al., 2023) B2C (General) No 1.41M 16.79M 1 No / No Yes No Yes Yes No Yes Yes ? 3.96M
DIGINETICA ? 232.93K 184.04K 3.3M 4 1217 / No No Yes No No Yes 941.64K ? ? 573.93K
YOOCHOOSE ? No 52.73K 34.15M 1 348 / No No Yes No No Yes No ? ? 9.24M
Retailrocket (Zykov et al., 2022) ? 1.4M 417.05K 2.75M 3 1669 / Yes No No No No Yes 51.29M ? ? No
Ali-CCP (Ma et al., 2018a) C2C (General) 400K 4.3M 87.41M 3 ? / ? Yes No No No No No ? ? No
Alibaba-iFashion (Chen et al., 2019) C2C (Fashion) 3.56M 4.46M 191.39M 1 75 / No No No No No No 7.7M ? ? No
@@ 4. Experiment & Analysis
In this section, we present the application of various machine learning and recommendation models to specific tasks using the MerRec dataset, demonstrating their effectiveness and performance outcomes. We refer to a previous work (Yuan et al., 2022) for the implementations of most of these models, with necessary modifications to accommodate the unique characteristics of the MerRec dataset.
@@ 4.1. Click-Through Rate (CTR) Prediction
To limit the total amount of computation time and resource, we chose to omit comprehensive hyperparameter search and reduce the sample used specifically for the benchmark executions. We use the first among six months of data in MerRec. We set the shortest input history window sequence in this execution to be 7, where we generate the binary prediction on the 8th item. Sequences equal to 8 events become one snapshot row and those longer than 8 events become multiple snapshot rows using the rolling window [^7] We chose this format in line with common practices in the CTR literature, e.g. (Shen, 2017; Yuan et al., 2022).. This implies we did not pad sequences shorter than 8 events (which does not exist in MerRec) to become longer sequences unlike TenRec (Yuan et al., 2022). Due to time constraints we did not specifically study the positive to negative sampling ratio on MerRec and how it affects training quality of the benchmark models, thus no negative sampling is enforced in this work and is left to the reader to explore.
@@ 4.2. Session-based Recommendation (SBR)
For Mercari’s C2C marketplace, each item is a single quantity listing, where it can be purchased only once and it needs to be removed from the predictions after being purchased. Additionally, with a large number of item listings created daily, the output space of the recommendation model will need to be updated constantly. For these reasons, existing language model based sequential recommenders are ill-suited for a dynamic C2C marketplace like Mercari.
@@
In response to the unique challenges of a Consumer-to-Consumer (C2C) marketplace, we developed Mercatran, a Three-Tower Transformer-based architecture. Inspired by the Two-Tower model used in YouTube recommendations by Covington et al. (Covington et al., 2016), Mercatran is tailored for dynamic marketplaces like Mercari to offer product-independent recommendations, leveraging user history to generate diverse and multistep recommendations.
Unlike the model proposed by Covington et al. (Covington et al., 2016) where items (videos) are represented by unique id’s, in Mercatran, we encode items with the content features only, similar to Wang et al. (Wang et al., 2021a). The output of the model is modified to produce trained embeddings directly instead of tokens (corresponding to words in a language model). These trained embeddings can then be used to perform similarity search in a vector database to map the embeddings to items.
@@
Both Covington et al. (Covington et al., 2016) and Wang et al. (Wang et al., 2021a) use a two-tower architecture to encode users and items, however, we found this architecture to be lacking in diversity where the recommendations generated are similar to the last interacted item. The model is unable to learn from older interactions even when using GRUs for the user tower (Wang et al., 2021a), since the output is a single embedding which can’t capture diverse user history. We built a solution conceptually similar to this called Mercatran V1 as the precursor of our current implementation. While it was very successful in driving orders, conversion and views (see Table 6), we observed lack of diversity, where the model would only exploit the most recent interaction to recommend items.
@@
To capture longer user-history and generate multi-time step recommendations, we propose a Three-Tower architecture called Mercatran V2, where we use Two-Towers to encodes users and a third tower to encode items. The two-towers to encode users are needed to generate recommendations into the future auto-regressively, where the first tower is a standard transformer-based encoder which can encode long user-history and the second tower is a transformer-based decoder that can generate recommendations into the future. A third tower (a transfomer-based encoder, we call item encoder) is needed to provide feedback to the user towers during training and it is fed the ground truth. To train the model, we optimize a symmetric cross-entropy loss over the cosine similarity scores of the predictions from the user-decoder and the associated predictions from the item-encoder, for each step, similar to Sohn (Sohn, 2016) and Radford et al. (Radford et al., 2021).
@@
At inference time, the user towers are used to generate multi-step embeddings that can be used to query items in a vector database, while the third tower is used to create embeddings to index items in the vector database. The third tower can be a standard MLP. We use a transformer-based encoder without self-attention to function as the third tower.
@@
Figure 9 shows the overall architecture of our three-tower model. In the current implementation, given user-history, at maximum 22 interactions, the model predicts the next 4 interactions (as embeddings), and these embeddings are used to perform a similarity search inside a vector database to find the recommended items. While this system does suffer from sampling bias (Yang et al., 2020; Yi et al., 2019) due to the training data only containing user interactions, generating multi-step recommendations introduces diversity in the results aiding in providing more visibility to a more diverse set of items.
To increase the overall model efficiency in terms of the effective time horizon covered by events included in the attention window, we deduplicated consecutive events in the dataset (mentioned in Section 3.2) if both the action type and item IDs remained the same per user, as we noticed that raw events without deduplication (e.g. buyers may repeatedly view the same item in consecutive events) often lead to a substantial reduction in sequential recommendation diversity and generally over-focusing on immediate-history trends. To further increase controllable diversity, we also force the model to generate multiple user encoding vectors when performing retrieval [^9] During online tests, we also found engagement improvements by reusing cached snapshots of older user encoding vectors, as well as using a reranker that promotes more randomness in ranking. This model-agnostic preprocessing is different from TransAct (Xia et al., 2023) which tackled a similar problem at Pinterest but used a mix of model-side methods such as masking a random-length block of immediate-history events during training.
@@
Table 5. Offline evaluation of Mercatran with 3 variations of input features, each measured at 3 levels of granularity, item (fine), category (coarse) and brand (coarse). The steps represent the 4 forecasted steps into the future. As can be seen Mercatran maintains strong numbers even at 4 steps into the future.
@@
Input Features Pred. Type Step nDCG @5 nDCG @20 Recall @5 Recall @20 Tr./Val Time (Hrs) VRAM (GB)
Title + Brand + Category Item 1 0.0407 0.0615 0.0789 0.1750 11.2 8.4
@@
Title Item 1 0.0410 0.0611 0.0802 0.1735 23.2 8.4
2 0.0402 0.0603 0.0780 0.1710
3 0.0402 0.0598 0.0786 0.1697
4 0.0399 0.0594 0.0771 0.1679
Category 1 N/A N/A 0.4524 0.5988
2 N/A N/A 0.4511 0.5981
3 N/A N/A 0.4503 0.5974
4 N/A N/A 0.4493 0.5975
Brand 1 N/A N/A 0.5024 0.6256
2 N/A N/A 0.5031 0.6268
3 N/A N/A 0.5038 0.6271
4 N/A N/A 0.5037 0.6264
Brand + Category Item 1 0.0111 0.0196 0.0217 0.0609 23.2 8.4
2 0.0111 0.0194 0.0215 0.0593
3 0.0113 0.0193 0.0217 0.0583
4 0.0110 0.0189 0.0211 0.0575
Category 1 N/A N/A 0.2769 0.3604
2 N/A N/A 0.2810 0.3626
3 N/A N/A 0.2818 0.3636
4 N/A N/A 0.2806 0.3638
Brand 1 N/A N/A 0.3020 0.3485
2 N/A N/A 0.3099 0.3587
3 N/A N/A 0.3125 0.3643
4 N/A N/A 0.3146 0.3680
4.2.7. Serving Mercatran.
Mercatran is served using an open-source serving framework on a public cloud model prediction service. The prediction service can scale up to many nodes, depending upon traffic, where each node has 1 T4 GPU and 8 vCPUs and 30GB of RAM. With 1 such node, a single online inference to generate embedding for 1 item can take 40 - 50 MS on average, while a single inference to generate user embeddings for 1 user (note the model generates 4 embeddings) can take 60 - 70 MS on average. In general, many such nodes are ran online with load balancers, caching and fault tolerance handling to take on the enormous production traffic with minimal downtime.
To reduce the number of times Mercatran needs to be invoked at the time of request, we pre-compute user and item embeddings as much as possible with 2 offline asynchronous batch-jobs.
  • As users interact with our system, we pre-calculate embeddings corresponding to their interactions with Mercatran’s User Towers, which are then stored in a feature store. Upon a user request, instead of invoking Mercatran directly, we first perform a lookup in the feature store to retrieve user embeddings.
  • As new items are added into the system, Mercatran’s Item Tower is used to embed these items in order to index them in a Vector Database. This is also performed in a batch setting
Note that these offline jobs are run close to real time with streaming in order to keep the recommendations fresh.
@@
For ablation study on input features, we experimented on Mercatran with combination subsets of item title, brand and category label text. We see models trained on item title and item title + brand + category out perform models trained on only brand + category. This highlights the importance of item level features over just high level brand and category identifiers, since users might be interested in the some small subset within a brand or category not fully represented or discriminated by the identifiers alone. We do see that a model trained with only item title outperforms the model trained on item title, brand and category in many metrics. Since the difference is small, we think this can be attributed to the stochasticity introduced by sampling (since it’s not possible to evaluate on the entire marketplace) and/or feature redundancy. Despite this our production model uses all three due to additional information provided by the brand and category features.